OpenStudy (anonymous):

Find the measures of the angles, given that x+y=40.(Hint: You must solve a system of equations.) (x+2y)+(11x)=180 Show work

OpenStudy (anonymous):

Is the 2nd equation this: (x+2y)+(11x)=180 or this: (x+2y)*(11x)=180

OpenStudy (anonymous):

No it's plus sign between them (x+2y)+(11x)=180

OpenStudy (anonymous):

1st equation: x+y=40 y=40-x 2nd equation: x+2y+11x=180 12x+2y=180 6x+y=90 Substituting y with (40-x) 6x+40-x=90 5x=50 x=10 y=40-x=40-10=30 So x=10 , y=30

OpenStudy (anonymous):

I understood everything except the last one y=40-x=40-10=30 So x=10, y=30 can u explain the last part more

OpenStudy (anonymous):

no never mind I got it thank u so much that was very helpful